# WonderPlay: Dynamic 3D Scene Generation from a Single Image and Actions

### v1

@@ 3 WonderPlay
Background. We represent the background with Fast Layered Gaussian Surfels (FLAGS) [72], which can be seen as a simplified version of 3D Gaussian Splatting [32] designed specifically for single-image 3D scene generation. Formally, the background $\mathcal{B}_{t}=\{\mathbf{p}^{\text{B}},\mathbf{q}^{\text{B}},\mathbf{s}^{\text{B}},\mathbf{o}^{\text{B}},\mathbf{c}_{t}^{\text{B}}\}$ consists of $N_{\text{B}}$ Gaussian surfels, parameterized by 3D spatial positions $\mathbf{p}^{\text{B}}\in\mathbb{R}^{3N_{\text{B}}}$, orientation quaternions $\mathbf{q}^{\text{B}}$, scales $\mathbf{s}^{\text{B}}$, opacities $\mathbf{o}^{\text{B}}$, and view-independent RGB colors $\mathbf{c}^{\text{B}}_{t}$. We only allow the color to change over time to account for shading effects. In a nutshell, the initial scene background $\mathcal{B}_{0}$ is generated by decomposing the input image $\mathbf{I}$ into several image layers, unprojecting all pixels in each layer to 3D space with estimated depth [31], followed by a photometric optimization to match the rendering with the input image $\mathbf{I}$ via differentiable rendering [32]. We refer the reader to Yu et al. [72] for more details of the generation process. In our physics solver, we treat the background as a static boundary.
@@
Materials. Besides the geometry and appearance representation $\mathcal{O}$, an object also has material properties $\mathbf{m}$. The definition of object material depends on the object type, which follows a 6-way classification: rigid, elastic, cloth, smoke, liquid, and granular. We detail the material properties in Appendix B.2. We consider homogeneous uniform materials, i.e., $\mathbf{m}$ is constant within an object. We follow Liu et al. [43] to estimate the values of the material parameters $\mathbf{m}$ by a Vision-Language Model (VLM) with optional manual adjustment for physical plausibility during simulation.
@@ 3.2 Hybrid Generative Simulator
where we keep all non-dynamics attributes the same as $\mathcal{S}_{0}$. To compute the dynamics attributes of various materials, we employ multiple types of physics solvers. These solvers are coupled to tackle multi-physics scenes, e.g., fluid and rigid as shown in Figure 2. We briefly describe our solver for each material below, and provide details in Appendix B.2.
Rigid body. We adopt a rigid body solver for scenes with rigid objects based on shape matching [47]. At each simulation time step, the rigid solver uses action forces to update the dynamics attributes, and then detects collisions among rigid objects and the background using connectivity information $\mathbf{E}$ to resolve penetrations.
Elastic, granular materials, and liquid. We model these effects with continuum mechanics and simulate them using a Material-Point-Method (MPM) solver [30]. At each time step, an MPM solver computes the momentum of each object particle (surfel) to update the dynamics attributes.
Smoke and cloth. We employ the Position-Based Dynamics (PBD) solver [48] for these effects. At each timestep, the PBD solver predicts new positions for the particles, and then corrects them by resolving constraint violations. The constraints for smoke include incompressibility [44]; the constraints for cloth include stretch and bending compliance.
@@
Discussion. With our bimodal control, we pass the coarse motion and appearance information from the physics simulator to the video generator. This allows not only generating more realistic motion, but also fixes appearance artifacts caused by imperfect 3D scene reconstruction and the lack of lighting information to resolve appearance changes. The question is how much we trust the generator to overwrite the coarse information. Intuitively, this is dictated by $s_{1}$: If $s_{1}$ is close to 00, then the generator only modifies the coarse video $\tilde{\mathbf{V}}$ a bit; if $s_{1}$ is close to $S$, then it can overwrite $\tilde{\mathbf{V}}$ more and hallucinate new contents. Thus, $s_{1}$ positively corresponds to the responsibility of the video generator.
@@
where $\mathbf{V}_{s_{2}}$ is computed from gradually denoising $\mathbf{V}_{s_{1}}$. $\mathbf{M}\in\{0,1\}^{(T+1)\times H\times W\times 3}$ denotes the binary mask that uses $1$ to mark a pixel of dynamic objects and 00 to mark a pixel of the background, which is rendered from the coarse scene. After this computation, we set $\mathbf{V}_{s_{2}}\leftarrow\hat{\mathbf{V}}_{s_{2}}$ and continue the denoising to generate $\mathbf{V}$.
@@ 4 Experiments
Implementation details. For physics simulation, we adopt the Genesis [3] framework, which unifies several different physics solvers. For all scenes, we run physical simulation for $960$ steps and render one frame for each $20$ steps. When conditioning the video generator with simulated dynamics, we use the standard resolution and time values: $H=480$, $W=720$, with $T=48$ frames (in total $49$ frames output). For sampling, we use a DDIM [57] scheduler and iterate $S=25$ steps on the warped noise for the final video output. We empirically set degradation factor $\gamma=0.4$ and apply appearance signal at $s_{1}=21,s_{2}=18$ diffusion steps , as this combination usually provides the optimal results. We include additional implementation details in Appendix B.3.
@@ 4.1 Results
Despite their ability to produce plausible visual quality, video generation methods struggle to adhere to the actions. In the duck-dropping-into-water scene, the drag-conditioned model Tora [74] submerges the duck under the water and then changes its shape after it re-emerges. CogVideoX-I2V struggles to generate realistic dynamics for the duck’s drop and adds undesirable dynamics by moving the duck to the left. Both models also struggle with the boats scene. Tora completely alters the scene mid-video, while CogVideoX-I2V fails to generate meaningful dynamics.
As for the physics-based method, PhysGen [43] is limited to rigid body simulation in 2D space, making it hard to handle scene with complex materials such as water. PhysGaussian [43] typically requires multiview images and as a result, struggles to produce a reasonable 3D representation with only one input image. Also due to the lack of a complete 3D physical state, both physics-based methods fail to properly handle the shading effect in the boats scene and make the reflections move with the boats. Our method, in contrast, offers the advantages of both physical simulation and video generation: the physical simulation handles a wide range of materials and ensures the desired dynamics, and the video generation model provides visual realism by successfully synthesizing water waves and bubbles surrounding the duck, as well as the following reflections.
The qualitative observations are backed by the quantitative results in Table 2. Compared to both physics-based methods and video models, WonderPlay (ours) achieve the best or second-best performance across all metrics, showing strong motion quality, visual quality, and physical plausibility.
@@
Diverse scenes and materials. In Figure 5, we present the generated dynamic 3D scenes on a variety of input images with diverse actions. It is important to note that achieving realistic visual quality in simulations of complex materials from a single image input with limited physical state information is extremely challenging. However, with the aid of the video generator, the sticky jam appears vivid as it pours onto the cake, and the river waves look natural in response to the boat’s movement. Notably, the underlying physical simulator ensures that all dynamics follow the input actions. For example, the roses are initially blown to the right by the wind and then move back due to their elasticity, and the force on the snowman’s head pushes it to the right while the lower part remains intact—dynamics that are difficult for existing video generation models to capture.
Condition on different actions. A significant advantage of our method is that it enables generating different interactions with different actions in the same scene. In Figure 6, we present four scenes, each with two different actions and their corresponding output dynamic 3D scenes. For the top-left scene, the shoe’s action is moving left and then right, with the smoke being appropriately affected by the shoe’s movement. In the bottom-right scene, we demonstrate different vortical wind fields, one rotating parallel to the image plane and the other rotating perpendicular to it, with the kites moving accordingly in response to the wind force.
@@ 5 Conclusion
Acknowledgements. We thank Guandao Yang, Yunzhi Zhang, and Zhehao Li for the comments and fruitful discussions, and Hadi Alzayer for help reviewing the draft. This work is in part supported by NSF RI #2211258 and #2338203, ONR YIP N00014-24-1-2117, and ONR MURI N00014-22-1-2740.
@@ Appendix B Technical Details
B.1 Reconstructing Topological Gaussian Surfels
@@
B.2 Material and Physics Solvers
In the main paper we briefly describe our material model for each category of objects and their solvers. Here we provide further complementary information on them. To model an object, our first step is to do a 6-way classification (rigid, elastic, cloth, smoke, liquid, and granular) by a VLM. We assume the material properties are constant for all particles in an object. The material models and solvers are as follows.
@@ B.2 Material and Physics Solvers
Rigid body. We model a rigid object as a strictly undeformable mesh without internal links. The material properties $\mathbf{m}$ of a rigid object includes the density $\rho$ and the friction coefficient $k$. Recall that our topological Gaussian surfels are given by: $\mathcal{O}_{t}=\{\mathbf{E},\mathbf{v}_{t},\mathbf{p}_{t}^{\text{O}},\mathbf{q}_{t}^{\text{O}},\mathbf{s}_{t}^{\text{O}},\mathbf{o}_{t}^{\text{O}},\mathbf{c}_{t}^{\text{O}}\}$, where the edge matrix $\mathbf{E}\in\{0,1\}^{N_{\text{O}}\times N_{\text{O}}}$ indicates the topological connectivity of the surfels, and $\mathbf{v}_{t}\in\mathbb{R}^{3N_{\text{O}}}$ denotes the velocity. They can be seen as a super-set of a mesh that has $\mathbf{E}$ and $\mathbf{p}$. Therefore, we can directly apply a rigid body solver to simulate our rigid objects.
Elastic, liquid, and granular materials. We model these materials with continuum mechanics, similar to PhysGaussian [67]. The material properties $\mathbf{m}$ include the density $\rho$, Young’s modulus $E$, and Poisson’s ratio $\nu$. For granular material, the material properties also include the friction angle $\theta$. The physics solver is built upon MPM [30], a hybrid Eulerian-Langrangian method. It simulates based on both particles and a spatial grid. As mentioned above, we densely sample particles inside the object in addition to the surface surfels. In each simulation time step, the momentum of each particle is transferred to the grid within a particle-to-grid step, to further compute the terms like deformation gradient. These updates are back propagated into particles through a grid-to-particle process to update particle dynamics properties like position and velocity.
Cloth and smoke. We model smoke and cloth with only particles. The material properties $\mathbf{m}$ for cloth includes density $\rho$ and stretch/bending compliance $p$. The material properties for smoke includes $\rho$ and viscosity coefficient $\mu$. The physics solver is Position-Based Dynamics (PBD) [48]. We also densely sample particles inside smoke. Unlike MPM method, PBD method directly models the positions of each particle through a list of inequality and non-equality constraints. In each time step, PBD solver solves each constraint sequentially and directly update the particle’s position which is then used to update the velocity. We refer the read to Bender et al. [9] for more information.
B.3 Simulation Parameters
@@
B.4 Rendering Simulated Dynamics

### v2

@@ 3 WonderPlay
Background. We represent the background with Fast Layered Gaussian Surfels (FLAGS) [72]. Formally, the background $\mathcal{B}_{t}=\{\mathbf{p}^{\text{B}},\mathbf{q}^{\text{B}},\mathbf{s}^{\text{B}},\mathbf{o}^{\text{B}},\mathbf{c}_{t}^{\text{B}}\}$ consists of $N_{\text{B}}$ Gaussian surfels, parameterized by 3D spatial positions $\mathbf{p}^{\text{B}}\in\mathbb{R}^{3N_{\text{B}}}$, orientation quaternions $\mathbf{q}^{\text{B}}$, scales $\mathbf{s}^{\text{B}}$, opacities $\mathbf{o}^{\text{B}}$, and view-independent RGB colors $\mathbf{c}^{\text{B}}_{t}$. We detail the reconstruction of background in Appendix B.2 and treat it as a static boundary in simulation.
@@
Materials. Besides the geometry and appearance representation $\mathcal{O}$, an object also has material properties $\mathbf{m}$. The definition of object material depends on the object type, which follows a 6-way classification: rigid, elastic, cloth, smoke, liquid, and granular. We detail the material properties in Appendix B.4.
@@ 3.2 Hybrid Generative Simulator
where we keep all non-dynamics attributes the same as $\mathcal{S}_{0}$. To compute the dynamics attributes of various materials, we employ multiple types of physics solvers. These solvers are coupled to tackle multi-physics scenes, e.g., fluid and rigid as shown in Figure 2. Please see details for different solvers in Appendix B.4.
@@
Discussion. With our bimodal control, we pass the coarse motion and appearance information from the physics simulator to the video generator. This allows not only generating more realistic motion, but also fixes appearance artifacts caused by imperfect 3D scene reconstruction and the lack of lighting information to resolve appearance changes. The question is how much we trust the generator to overwrite the coarse information. Intuitively, this is dictated by $s_{1}$: If $s_{1}$ is close to 0, then the generator only modifies the coarse video $\tilde{\mathbf{V}}$ a bit; if $s_{1}$ is close to $S$, then it can overwrite $\tilde{\mathbf{V}}$ more and hallucinate new contents. Thus, $s_{1}$ positively corresponds to the responsibility of the video generator.
@@
where $\mathbf{V}_{s_{2}}$ is computed from gradually denoising $\mathbf{V}_{s_{1}}$. $\mathbf{M}\in\{0,1\}^{(T+1)\times H\times W\times 3}$ denotes the binary mask that uses $1$ to mark a pixel of dynamic objects and 0 to mark a pixel of the background, which is rendered from the coarse scene. After this computation, we set $\mathbf{V}_{s_{2}}\leftarrow\hat{\mathbf{V}}_{s_{2}}$ and continue the denoising to generate $\mathbf{V}$.
@@ 4 Experiments
Implementation details. For physics simulation, we adopt the Genesis [3] framework, which unifies several different physics solvers. For all scenes, we run physical simulation for $960$ steps and render one frame for each $20$ steps. We include additional implementation details in Appendix B.1.
@@ 4.1 Results
Despite their ability to produce plausible visual quality, video generation methods struggle to adhere to the actions. In the duck-dropping-into-water scene, Tora [74] submerges the duck under the water and then changes its shape after it re-emerges. CogVideoX-I2V struggles to generate realistic dynamics for the duck’s drop and adds undesirable dynamics by moving the duck to the left. Both models also struggle with the boats scene. Tora completely alters the scene mid-video, while CogVideoX-I2V fails to generate meaningful dynamics.
As for the physics-based method, PhysGen [43] is limited to rigid body simulation in 2D space, making it hard to handle scene with complex materials such as water. PhysGaussian [43] typically requires multiview images and as a result, struggles to produce a reasonable 3D representation with only one input image. Also due to the lack of a complete 3D physical state, both physics-based methods fail to properly handle the shading effect in the boats scene and make the reflections move with the boats. Our method, in contrast, offers the advantages of both physical simulation and video generation: the physical simulation handles a wide range of materials and ensures the desired dynamics, and the video generation model provides visual realism by successfully synthesizing water waves and bubbles surrounding the duck, as well as the following reflections. Also shown in Table 2, WonderPlay (ours) achieve the best or second-best performance across all metrics, showing strong motion quality, visual quality, and physical plausibility.
@@
Diverse scenes and materials. In Figure 5, we present the generated dynamic 3D scenes on a variety of input images with diverse actions. It is important to note that achieving realistic visual quality in simulations of complex materials from a single image input with limited physical state information is extremely challenging. However, with the aid of the video generator, the sticky jam appears vivid as it pours onto the cake, and the river waves look natural in response to the boat’s movement. Notably, the underlying physical simulator ensures that all dynamics follow the input actions. For example, the roses are initially blown to the right by the wind and then move back due to their elasticity.
Condition on different actions. A significant advantage of our method is that it enables generating different interactions with different actions in the same scene. In Figure 6, we present four scenes, each with two different actions and their corresponding output dynamic 3D scenes.
@@ 5 Conclusion
Acknowledgments. We thank Guandao Yang, Yunzhi Zhang, and Zhehao Li for the comments and fruitful discussions, and Hadi Alzayer for help reviewing the draft. This work is in part supported by the Stanford Institute for Human-Centered AI (HAI), the Okawa Foundation Research Grant, NSF RI #2211258 and #2338203, ONR YIP N00014-24-1-2117, and ONR MURI N00014-22-1-2740.
@@ Appendix B Technical Details
B.1 Additional Implementation Details
When conditioning the video generator with simulated dynamics, we use the standard resolution and time values: $H=480$, $W=720$, with $T=48$ frames (in total $49$ frames output). For sampling, we use a DDIM [57] scheduler and iterate $S=25$ steps on the warped noise for the final video output. We empirically set degradation factor $\gamma=0.4$ and apply appearance signal at $s_{1}=21,s_{2}=18$ diffusion steps , as this combination usually provides the optimal results.
B.2 Reconstructing Background
In a nutshell, the initial scene background $\mathcal{B}_{0}$ is generated by decomposing the input image $\mathbf{I}$ into several image layers, unprojecting all pixels in each layer to 3D space with estimated depth [31], followed by a photometric optimization to match the rendering with the input image $\mathbf{I}$ via differentiable rendering [32]. We refer the reader to Yu et al. [72] for more details of the generation process.
B.3 Reconstructing Topological Gaussian Surfels
@@
B.4 Material and Physics Solvers
We consider homogeneous uniform materials, i.e., $\mathbf{m}$ is constant within an object. We follow Liu et al. [43] to estimate the values of the material parameters $\mathbf{m}$ by a Vision-Language Model (VLM) with optional manual adjustment for physical plausibility during simulation.
Here we provide further complementary information on each object material model and their solvers. We consider homogeneous uniform materials, i.e., $\mathbf{m}$ is constant within an object. To model an object, we follow Liu et al. [43] to do a 6-way classification (rigid, elastic, cloth, smoke, liquid, and granular) by a VLM, and estimate the values of the material parameters $\mathbf{m}$ by the VLM with optional manual adjustment for physical plausibility during simulation. The material models and solvers are as follows.
@@ B.4 Material and Physics Solvers
Rigid body. We model a rigid object as a strictly undeformable mesh without internal links. The material properties $\mathbf{m}$ of a rigid object includes the density $\rho$ and the friction coefficient $k$. Recall that our topological Gaussian surfels are given by: $\mathcal{O}_{t}=\{\mathbf{E},\mathbf{v}_{t},\mathbf{p}_{t}^{\text{O}},\mathbf{q}_{t}^{\text{O}},\mathbf{s}_{t}^{\text{O}},\mathbf{o}_{t}^{\text{O}},\mathbf{c}_{t}^{\text{O}}\}$, where the edge matrix $\mathbf{E}\in\{0,1\}^{N_{\text{O}}\times N_{\text{O}}}$ indicates the topological connectivity of the surfels, and $\mathbf{v}_{t}\in\mathbb{R}^{3N_{\text{O}}}$ denotes the velocity. They can be seen as a super-set of a mesh that has $\mathbf{E}$ and $\mathbf{p}$. Therefore, we can directly apply a rigid body solver to simulate our rigid objects. We adopt a rigid body solver based on shape matching [47]. At each simulation time step, the rigid solver uses action forces to update the dynamics attributes, and then detects collisions among rigid objects and the background using connectivity information $\mathbf{E}$ to resolve penetrations.
Elastic, liquid, and granular materials. We model these materials with continuum mechanics and simulate them using a Material-Point-Method (MPM) solver [30], similar to PhysGaussian [67]. The material properties $\mathbf{m}$ include the density $\rho$, Young’s modulus $E$, and Poisson’s ratio $\nu$. For granular material, the material properties also include the friction angle $\theta$. The physics solver is built upon MPM [30], a hybrid Eulerian-Langrangian method. It simulates based on both particles and a spatial grid. As mentioned above, we densely sample particles inside the object in addition to the surface surfels. In each simulation time step, an MPM solver computes the momentum of each object particle (surfel) to update the dynamics attributes. In detail, the momentum of each particle is transferred to the grid within a particle-to-grid step, to further compute the terms like deformation gradient. These updates are back propagated into particles through a grid-to-particle process to update particle dynamics properties like position and velocity.
Cloth and smoke. We model smoke and cloth with only particles and employ the Position-Based Dynamics (PBD) solver [48] for these effects. The material properties $\mathbf{m}$ for cloth includes density $\rho$ and stretch/bending compliance $p$. The material properties for smoke includes $\rho$ and viscosity coefficient $\mu$. We also densely sample particles inside smoke. Unlike MPM method, PBD method directly models the positions of each particle through a list of inequality and non-equality constraints. In each time step, PBD solver solves each constraint sequentially and directly update the particle’s position which is then used to update the velocity. The constraints for smoke include incompressibility [44]; the constraints for cloth include stretch and bending compliance. We refer the read to Bender et al. [9] for more information.
B.5 Simulation Parameters
@@
B.6 Rendering Simulated Dynamics